OpenStudy (anonymous):

what is the 2nd derivative of y^2=10x

OpenStudy (zarkon):

can you find the 1st derivative?

OpenStudy (anonymous):

yeah the first is 10/2y

OpenStudy (zarkon):

so \[y'=\frac{5}{y}\] take the derivative again

OpenStudy (anonymous):

I got \[-200/2y^3\] can you show me the work you did to get your answer

OpenStudy (zarkon):

I don't get that as an aswer

OpenStudy (zarkon):

\[y''=5\frac{-1}{y^2}y'=-5\frac{1}{y^2}\frac{5}{y}=-\frac{25}{y^3}\]

OpenStudy (anonymous):

i got it now thanks a lot

OpenStudy (zarkon):

no problem